OpenStudy (anonymous):

Solve for y. 3/y + 7/y = 9 A. 10/9 B. 5/9 C.90 I think the answer is A but not sure can anyone help and make sure I'm right

OpenStudy (anonymous):

a

jimthompson5910 (jim_thompson5910):

3/y + 7/y = 9 (3 + 7)/y = 9 10/y = 9 10 = 9y 9y = 10 y = 10/9 I'll let you check the answer

OpenStudy (anonymous):

Thx